4. Find the complete solution to the system 2y z t = 1 2x - y - z -t = 3 3x 2z - 2t = -2
Answer #1

1 + 24-2 =1 2 x-y-z -t = 3 3x + 7+22-2t = -2 21 - -- ( 3 111 Consider augmented matrix (22 2 - 1 - ) Reduce Augmented matin t- د) arturitni له- لم - - 1 + + t = t - =( +6 : -1-
Answer #2

Hello.

Please someone should help with the matlab code. 

thanks 

answered by: Anumah, Adeiza Stephen